Homework Question 1: Find the quantity and price that maximize profit for a firm with demand Q 133.33 1/3)P and cost function is C(Q) 25000 Q2 Answer to Homework Question 1
Answer #1

Profit is maximized where MR = MC

P = 400 - 3Q

TR = PQ = 400Q - 3Q2

MR = d(TR)/dQ = 400 - 6Q

MC = dC/dQ = 2Q

Equating MR with MC:

400 = 8Q

Q = 50 and P = 400 - 3 x 50 = 250